OpenStudy (anonymous):

Can someone help me with a worksheet on Fundamental Theorem of Algebra and Descaretes rule of signs

OpenStudy (anonymous):

got a question?

OpenStudy (anonymous):

Its the same as yesterday just a different function.

OpenStudy (anonymous):

my memory is not that good you got a specific function ?`

OpenStudy (anonymous):

Use the Fundamental Theorem of Algebra to determine the total number of zeros of f(x)=x^4-x^3+7x^2-9x-18

OpenStudy (anonymous):

it is a polynomial of degree 4 there are 4 zeros although some might be complex and some might be repeated do you need to find possible number of positive and negative zeros, or do you just need to find the actual zeros?

OpenStudy (anonymous):

i need help on a whole worksheet on this I know most of it just a few questions i need help with

OpenStudy (anonymous):

i got ten minutes and can help if you post a screenshot

OpenStudy (anonymous):

i have 1-3c done and i know how to do 4a and b

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

lets go to descartes rule of sign check the changes in sign of the coefficient of \(x^4-x^3+7x^2-9x-18\)

OpenStudy (anonymous):

there are 3 from \(1\) to \(-1\) from \(-1\) to \(7\) and from \(7\) to \(-9\_

OpenStudy (anonymous):

that means there are either 3 positive zeros ,or one positive zero (you count down by twos)

OpenStudy (anonymous):

now given \[f(x)=x^4-x^3+7x^2-9x-18\] compute \[f(-x)=x^4+x^3+7x^2+9x-18\] and here there is only one change in sign from \(+9\) to \(-18\) that means there is a negative zeros

OpenStudy (anonymous):

gotta run, good luck

OpenStudy (anonymous):

So there would be 3?

OpenStudy (anonymous):

Alright thank you soo much for helping